Q2.Calculate Net Domestic Income, Gross Domestic


Income. Net national Income

ItemsRs Crores

Compensation of employees24000

Operating Surplus10000

Mixed Income of Self-employees28000

Net factor income earned from abroad-3000

Depreciation1700

Indirect Taxes9000

Subsidy1800


Ans:

Net domestic Income= 62000 Crores

Gross Domestic Income= 63700 Crores


Net National Income= 61700 Crores


1
Expert's answer
2021-07-06T18:21:10-0400

Gross domestic Income = compensation of employees + operating surplus + Mixed income of self employees + Depriciation

"= 24000 + 10000 + 28000 + 1700"

"= 63700"

Net domestic income = Gross domestic income - Depriciation

"= 63700 - 1700\\

= 62000"